0ryz. we 13.990.09.y 0.314+0.001 and Given the equation f -1 z 0.0580+0.0006 What is the absolute uncertainty in fr Select one: o -2.55 0.02 -2.6 0.2 o. 2.6+0.5 d. 2.55+0.05 Check
Answer #1

Result : (d) 2.55 pm0.05

Explanation:

  • Given: f = 10xyz , x = 13.99pm0.09, y= 0.314pm0.001 , z = 0.0580pm​​​​​​​0.0006

x*y*z = (13.99*0.314*0.0580)pmxyz[0.09 0.001 0.0006 13.99 0.314 0.0580]

= 0.255 pm 0.25(0.02)

= 0.255 pm 0.005

Now , f = 10xyz

So absolute uncertainty in f is given by:

f = (0.25×10)pm(0.005×10)

= 2.55 pm0.05